OpenStudy (barrelracing):

I really need help with this stuff but please explain, no direct answers write and equation in standard form the line has the same slope as 5x-y=7 and the same y-intercept as the graph of 2y-7x=8. the equation in standard form is?

OpenStudy (sdfgsdfgs):

eqn of a straight line in standard form is y = mx + c; where m is the slope and c is the y-intercept. So far so good?

OpenStudy (barrelracing):

ok

OpenStudy (sdfgsdfgs):

"he line has the same slope as 5x-y=7 " so we need to find out the slope from 5x-y=7.... can u put 5x-y=7 into the standard form of y=mx+c plz?

OpenStudy (barrelracing):

y=-5x+7

OpenStudy (barrelracing):

are you there?

OpenStudy (sdfgsdfgs):

Ooops sorry I got distracted....almost right but it should be y=5x-7 see where u made the mistake?

OpenStudy (barrelracing):

yes

OpenStudy (sdfgsdfgs):

So what is the slope? We need that for part of the ans :)

OpenStudy (barrelracing):

5

OpenStudy (sdfgsdfgs):

Correct! So keep that in mind and we work on finding the y-intercept ok?

OpenStudy (sdfgsdfgs):

"(T)he same y-intercept as the graph of 2y-7x=8" -same deal: put 2y-7x=8 into standard from y=mx+c to find the y-intercept

OpenStudy (barrelracing):

2y=7x+8 y=7/2x+4

OpenStudy (sdfgsdfgs):

GREAT job! :) so whats the y-intercept here?

OpenStudy (barrelracing):

4

OpenStudy (sdfgsdfgs):

Combining that w slope of 5, ur ans in standard form eqn is....

OpenStudy (barrelracing):

y=5x+4?

OpenStudy (barrelracing):

is that right?

OpenStudy (sdfgsdfgs):

Bingo - u got it :)

OpenStudy (barrelracing):

thank you sooo much!

OpenStudy (sdfgsdfgs):

welcome! U did most of the work urself! lol